OpenStudy (anonymous):

Solve A = 2πr + 3t for t.

OpenStudy (jhannybean):

We have \[\large A = 2 \pi r +3t \] We want to isolate the t to one side and everything else to the other side. In order to do this we start by subtracting 2 pi r from both sides. \[\large A - 2 \pi r = -2 \pi r +2 \pi r + 3t \] Then we simplify the equation. \[\large A - 2 \pi r = 3t \] now that we've got our t on one side and everything else on the otehr, we notice the "t" has a coefficient in front of it that we need to get rid of,we an divide both sides of the equation by 3 to get rid of it. \[\large \frac{A-2\pi r}{3}= \frac{3t}{3}\] Simplifying the equation again leaves us with \[\large \frac{A-2 \pi r}{3} = t\]

OpenStudy (anonymous):

thank you

OpenStudy (jhannybean):

Understand how its done? :P